OpenStudy (cenaida):

|2- 3x| = 7 A.{-5/3, 3} B.{-15, 3} C.{-3, 5/3} D.{-3, 15}

jimthompson5910 (jim_thompson5910):

same idea as last time |2-3x| = 7 turns into 2-3x = 7 or 2-3x = -7 solve each for x

OpenStudy (anonymous):

yup thats true

OpenStudy (cenaida):

but 3 cant go into 7

jimthompson5910 (jim_thompson5910):

2 - 3x = 7 -3x = 7 - 2 -3x = 5 x = ??

jimthompson5910 (jim_thompson5910):

that will give you one solution

OpenStudy (cenaida):

do i multiply it then? then it will be -15

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

you divide both sides by -3

jimthompson5910 (jim_thompson5910):

to isolate x

OpenStudy (cenaida):

but 3 cant go into 5 evenly

jimthompson5910 (jim_thompson5910):

that's why you'll leave it as a fraction

jimthompson5910 (jim_thompson5910):

or get a decimal approximation but in this case, they want the fraction

OpenStudy (cenaida):

oooh so its -5\3

jimthompson5910 (jim_thompson5910):

-5/3 yes

jimthompson5910 (jim_thompson5910):

that's one of them

OpenStudy (cenaida):

and the other one is 3 right?

jimthompson5910 (jim_thompson5910):

yes, since 2 - 3x = -7 -3x = -7 - 2 -3x = -9 x = -9/(-3) x = 3